# Exact scattering cross section for lattice-defect scattering of phonons

### v1

@@ Abstract
The use of structurally complex lattice defects, such as functional groups, embedded nanoparticles, and nanopillars, to generate phonon scattering is a popular approach in phonon engineering for thermoelectric applications. However, the theoretical treatment of this scattering phenomenon remains a formidable challenge, especially with regards to the determination of the scattering cross sections and rates associated with such lattice defects. Using the extended Atomistic Green’s Function (AGF) method, we describe how the numerically exact mode-resolved scattering cross section $\sigma$ can be computed for a phonon scattered by a single lattice defect. We illustrate the generality and utility of the AGF-based treatment with two examples. In the first example, we treat the isotopic scattering of phonons in a harmonic chain of atoms . In the second example, we treat the more complex problem of phonon scattering in a carbon nanotube (CNT) containing an encapsulated C60 molecule which acts as a scatterer of the CNT phonons. The application of this method can enable a more precise characterization of lattice-defect scattering and result in the more controlled use of nanostructuring and lattice defects in phonon engineering for thermoelectric applications.
@@ I Introduction
In this work, we propose an alternative approach for computing the scattering cross section $\sigma$, based on the extended Atomistic Green’s Function (AGF) method, (Ong and Zhang, 2015; Ong, 2018a, b) which is used commonly to study ballistic phonon transport (Mingo and Yang, 2003; Wang et al., 2006; Zhang et al., 2007a, b; Wang et al., 2008; Ong and Zhang, 2015) and has also been applied to model elastic wave scattering (Khodavirdi et al., 2024) as well as to characterize the boundary roughness scattering of phonons. (Ong et al., 2020; Ong, 2024) In the extended AGF method, which is a development of the original AGF method introduced by Mingo and Yang for treating the total phonon transmission in nanowires, (Mingo and Yang, 2003) mode-resolved quantities, such as the scattering amplitudes as well as mode-resolved transmission and reflection coefficients, can be calculated from the scattering $S$ matrix. (Ong, 2018b) The method for $\sigma$ is based on a rigorous interpretation of the scattering amplitude and thus does not require any new computational technique as $\sigma$ is derived from the diagonal element of the transmission matrix computed within the extended AGF method. (Ong, 2018b) For the sake of simplicity, we limit the discussion of our proposed approach to quasi-one-dimensional systems, such as nanotubes or nanowires, although the approach can be easily extended to two or three-dimensional systems, which we briefly discuss in Appendix B.
@@
Figure 1: Schematic of layered system in the AGF method. The inhomogeneity or defect, which is represented by the red-shaded oval, is contained within the layers $n=1$ to $N-1$ which correspond to the “Device” region. The distance between the left and right leads is $W=Na$.
@@ II Theoretical basis of AGF-based approach
In the AGF method, (Ong, 2018a, b) which is outlined in Appendix A, we arrange the degrees of freedom in an infinite array of layers, which we enumerate as $n=-\infty,\ldots,0,1,\ldots,\infty$ and assume are evenly spaced with an interlayer distance of $a$, as shown in Fig. 1. Thus, we can associate a discrete coordinate $x=na$ with the center of the $n$-th layer. As before, we partition the system into three subsystems: the left lead ($n\leq 0$), the device ($11<n<N-$ 1) and the right lead ($n\geq N$). In the left and right leads which are homogeneous and identical to one another, each layer acts as a unit cell, has $l$ degrees of freedom and is identical to the other cells in the lead. In the device, the layers are not necessarily homogeneous even though they are spaced evenly and can accommodate the additional degrees of freedom associated with an inhomogeneity. In our derivation, we discuss the scattering cross section associated with the inhomogeneity for an incoming phonon propagating rightward from the left lead towards the device.
@@ II.1 Total outgoing phonon flux from scattering
where $\bm{\Psi}_{\text{inc}}^{\text{L}}(x)=\bm{\Phi}_{\nu k}(x)\equiv\sqrt{|v_{\nu k}|}\bm{u}_{\nu k}(x)e^{ikx}$ represents the incident $\nu k$ phonon mode  [^2] A careful distinction should be made between the terms “incident” and “incoming”. The former is used to refer to the unperturbed state when the scatterer is absent while the latter is used to refer to the phonon mode that is traveling towards the device region. A rightward-propagating incident phonon consists of an incoming phonon eigenmode in the left lead and an outgoing phonon eigenmode in the right lead. with group velocity $v_{\nu k}$ and $\bm{u}_{\nu k}(x)$ denotes its Bloch function satisfying the relationship $\bm{u}_{\nu k}(x)=\bm{u}_{\nu k}(x+a)$. Because of flux normalization, Eq. (1) only contains the contribution from the propagating modes and none from the evanescent modes. In the absence of scattering, the phonon wave function at any two layers at $x$ and $x^{\prime}$ only differ by a phase factor determined by its wave vector, i.e., $\bm{\Psi}_{\text{L}}(x)=\bm{\Psi}_{\text{L}}(x^{\prime})e^{ik(x-x^{\prime})}$.
@@
which is written as a sum of the reflected leftward-propagating flux-normalized $\nu^{\prime}k^{\prime}$ eigenmodes associated with the scattering amplitudes $S(\nu^{\prime}k^{\prime},\nu k)$. [^3] The Heaviside term $\Theta(-v_{k^{\prime}})$ eliminates the contributions from the transmitted rightward-propagating waves in Eq. (2). The sum over $\nu^{\prime}k^{\prime}$ in Eq. (2) is over all the possible branches and wave vectors at the incident phonon frequency $\omega_{\nu k}$, i.e., $\omega_{\nu^{\prime}k^{\prime}}=\omega_{\nu k}$. We can associate two fluxes with the left-lead wave function: one from the incident mode and the other from the modes reflected from the device into the left lead. The normalized incident flux is $j_{\text{inc}}^{\text{L}}(\nu k)=1$ while the reflected flux is
@@
The RHS of Eq. (4) contains the incident and outgoing forward-scattered wave functions. Each summand has a phase factor of $e^{-ik^{\prime}W}$, where $W$ denotes the width of the scatterer and is taken as the distance between the lead edges (from $n=0$ in the left lead to $n=N$ in the right lead) so that we have $W=Na$. In the absence of a scatterer (e.g. homogeneous waveguide), the forward-scattering amplitude in the right lead is $\overline{S}(\nu^{\prime}k^{\prime},\nu k)=\delta_{\nu^{\prime}\nu}\delta_{k^{\prime},k}e^{ik^{\prime}W}$ where $\delta_{k^{\prime},k}$ denotes the Kronecker delta function, and we recover the solution $\overline{\bm{\Psi}}_{\text{R}}(x)=\sum_{\nu^{\prime}k^{\prime}}\overline{S}(\nu^{\prime}k^{\prime},\nu k)\bm{\Phi}_{\nu^{\prime}k^{\prime}}e^{-ik^{\prime}W}\Theta(v_{\nu^{\prime}k^{\prime}})=\bm{\Phi}_{\nu k}(x)$, which is just the incident eigenmode.
@@ II.2 Scattering cross section from the transmission-matrix diagonal element
In the absence of a scatterer, Eq. (9) yields $\sigma=0$ as expected since $S(\nu k,\nu k)=\overline{S}(\nu k,\nu k)$ and $\mathcal{F}(\nu k)=0$. On the other hand, we note that $\sigma\leq 4A$, i.e., the maximum value of the total scattering cross section can be as large as four times of the geometrical cross section when $S(\nu k,\nu k)=-\overline{S}(\nu k,\nu k)$ and $\mathcal{F}(\nu k)=-2$(Boya, 2008) While it appears to be counter-intuitive and even possibly unphysical, this result also occurs in quantum mechanics for scattering by a hard sphere in three dimensions. (Sakurai and Napolitano, 2020) This is a consequence of our partitioning of the wave function in $\bm{\Psi}_{\text{L}}(x)$ and $\bm{\Psi}_{\text{R}}(x)$ into two non-orthogonal components: the incident wave $\bm{\Psi}_{\text{inc}}$ and the scattered wave $\bm{\Psi}_{\text{scatt}}$ which represents the part of the original wave that has been perturbed. To find $S(\nu k,\nu k)$ for the $\nu k$ mode in the left lead, we use the AGF method to compute the $l\times l$ transmission matrix for the inhomogeneous waveguide, given by (Ong, 2018a, b)
@@
where $\omega=\omega_{\nu k}$ and $\boldsymbol{G}_{\text{RL}}^{\text{ret}}$ denotes the Green’s function between the rightmost left-lead ($n=0$) and the leftmost right-lead ($n=N$) layers. In Eq. (10),$\boldsymbol{V}_{\text{R}}^{\text{ret}}(+)$ and $\boldsymbol{V}_{\text{L}}^{\text{adv}}(-)$ correspond to the velocity matrices for the outgoing right-lead phonons and the incoming left-lead phonons, respectively, while $\boldsymbol{U}_{\text{R}}^{\text{ret}}(+)$ and $\boldsymbol{U}_{\text{L}}^{\text{adv}}(-)$ correspond to the outgoing right-lead eigenmodes and the incoming left-lead eigenmodes, respectively. A more detailed description of these matrices is given in Eqs. (20b) to (22) in Appendix A. The transmission matrix element $[\boldsymbol{t}_{\text{RL}}(\omega)]_{pq}$ connects the flux amplitude of the $p$-th outgoing right-lead modes to the $q$-th incoming left-lead modes for $p,q=1,\ldots,l$. Because the left and right leads are identical, the set of incoming extended left-lead modes, as indexed by $\nu k$, is the same as the set of outgoing extended right-lead modes, and we can order the extended phonon modes by their wave vectors such that $k_{1}<k_{2}<\ldots$.
@@ III.1 Monoatomic harmonic chain with a single isotopic impurity
To illustrate the AGF method for calculating the scattering cross section, we apply it to the simple archetypal example (Allen and Nghiem, 2022) of phonon scattering by a single isotopic impurity of mass $m_{\text{C}}$ in a monoatomic harmonic chain with an interatomic distance of $a$, as shown in Fig. 2(a). In this system, we treat the isotopic impurity as the “device”; the mass and interatomic force constants of the rest of the system are given by $m$ and $\zeta$, respectively. For the monoatomic harmonic chain, the phonon dispersion is given by $\omega(k)^{2}=\omega_{m}^{2}\sin^{2}(ka/2)$, where $\omega_{m}=2\sqrt{\zeta/m}$, and the phonon group velocity is given by $v(\omega)=\pm\frac{1}{2}a(\omega_{m}^{2}-\omega^{2})^{1/2}$. As there is only one phonon branch, all of the terms in Eq. (10) are scalars. Thus, Eq. (10) simplifies to $t_{\text{RL}}(\omega)=i\omega(\omega_{m}^{2}-\omega^{2})^{1/2}G_{\text{RL}}^{\text{ret}}(\omega)$.
@@ III.1.2 Scattering cross section and rate
After some algebra, we obtain from Eq. (12) the expression for the scattering cross section of the single isotopic impurity, normalized by the geometrical cross section $A$:
@@ III.2 Carbon nanotube with an encapsulated C60 molecule
Figure 3: Schematic of the layer arrangement in the (10,10) armchair carbon nanotube with a C60 molecule (pink-shaded atoms) encapsulated within the “Device” region (turqoise-shaded atoms) in layers $n=1$ to $12$. The layers in the leads, which extend semi-infinitely to the left or right, are represented by the yellow-shaded atoms.
@@
Figure 4: (a) Plot of scattering cross section $\sigma/A$ for each rightward-propagating phonon mode (solid circle) in the (10,10) armchair CNT, distributed over the 66 distinct phonon branches. The values of $\sigma/A$ are between 0 and 4 and are indicated by color according to the color bar. (b) The corresponding plot at low frequencies ($\omega<4\times 10^{13}$ rad/s). We observe the three distinct acoustic phonon branches (LA, TW and TA) and the radial breathing mode (BR) as well as the five lowest phonon subbands which are indicated by brown dashed arrows and labeled SB1 to SB5. The phonon branches are also colored according to the percentage of their eigenmode atomic displacements parallel to the longitudinal axis, with the color varying continuouly between bright magenta (100 percent) and black (0 percent).
@@ III.2.1 Distribution of mode-dependent scattering cross sections
We compute the area-normalized scattering cross section $\sigma(\nu k)/A$ for each rightward-propagating $\nu k$ phonon mode in the (10,10) armchair CNT using Eqs. (10) and (12) for a range of angular frequencies $\omega$ in intervals of $\Delta\omega=2\times 10^{12}$ rad/s from $\omega=2\times 10^{12}$ to $3.16\times 10^{14}$ rad/s. At each $\omega$, we obtain a set of phonon modes with the wave vectors $k_{1},k_{2},\ldots$ which we plot as solid circles in Fig. 4. Like in Fig. 2, we indicate the value of $\sigma/A$ of each mode by the color of its solid circle symbol. For ease of interpretation, the solid circle symbols are superimposed on the $\omega-k$ phonon dispersion curves for the (10,10) armchair CNT. (Dresselhaus and Eklund, 2000)
In Fig. 4(a), there are 66 distinct ones due to the degeneracies in the (10,10) armchair CNT. (Dresselhaus and Eklund, 2000) At $k=0$ in the $\omega\rightarrow 0$ limit, there are 4 acoustic phonon branches corresponding to the longitudinal acoustic (LA), the azimuthal twisting (TW) and the doubly degenerate radial transverse acoustic (TA) phonons. (Mahan, 2002; Chico et al., 2006) In addition, there is a breathing mode (BR) associated with the radial oscillation of the CNT at $k=0$. The remaining phonon branches consist of subbands with $\omega>0$ at $k=0$. We observe two distinct $\sigma/A$ trends for the lower-frequency phonon subbands in Fig. 4(a). In the subbands with higher group velocities and near-linear dispersion, $\sigma/A$ is vanishingly small because the atomic motion for the phonon modes is primarily along the longitudinal direction (Dresselhaus and Eklund, 2000). On the other hand, in the subbands with smaller group velocities and a more parabolic dispersion, $\sigma/A$ increases significantly with $\omega$, indicating stronger interaction with the C60 molecule. For the subbands correspond to the high-frequency optical phonons, we have $\sigma/A\sim 2$, which indicates strong scattering.

### v2

@@ Abstract
The use of structurally complex lattice defects, such as functional groups, embedded nanoparticles, and nanopillars, to generate phonon scattering is a popular approach in phonon engineering for thermoelectric applications. However, the theoretical treatment of this scattering phenomenon remains a formidable challenge, especially with regard to the determination of the scattering cross sections and rates associated with such lattice defects. Using the extended atomistic Green’s function (AGF) method, we describe how the numerically exact mode-resolved scattering cross section $\sigma$ can be computed for a phonon scattered by a single lattice defect. We illustrate the generality and utility of the AGF-based treatment with two examples. In the first example, we treat the isotopic scattering of phonons in a harmonic chain of atoms . In the second example, we treat the more complex problem of phonon scattering in a carbon nanotube (CNT) containing an encapsulated C60 molecule which acts as a scatterer of the CNT phonons. The application of this method can enable a more precise characterization of lattice-defect scattering and result in a more controlled use of nanostructuring and lattice defects in phonon engineering for thermoelectric applications.
@@ I Introduction
In this paper, we propose an alternative approach for computing the scattering cross section $\sigma$, based on the extended atomistic Green’s function (AGF) method, (Ong and Zhang, 2015; Ong, 2018a, b) which is commonly used to study ballistic phonon transport (Mingo and Yang, 2003; Wang et al., 2006; Zhang et al., 2007a, b; Wang et al., 2008; Ong and Zhang, 2015) and has also been applied to model elastic wave scattering (Khodavirdi et al., 2024) as well as to characterize the boundary roughness scattering of phonons. (Ong et al., 2020; Ong, 2024) In the extended AGF method, which is a development of the original AGF method introduced by Mingo and Yang for treating the total phonon transmission in nanowires, (Mingo and Yang, 2003) mode-resolved quantities, such as the scattering amplitudes as well as mode-resolved transmission and reflection coefficients, can be calculated from the scattering $S$ matrix. (Ong, 2018b) The method for $\sigma$ is based on a rigorous interpretation of the scattering amplitude and thus does not require any new computational technique as $\sigma$ is derived from the diagonal element of the transmission matrix computed within the extended AGF method. (Ong, 2018b) For the sake of simplicity, we limit the discussion of our proposed approach to quasi-one-dimensional systems, such as nanotubes or nanowires, although the approach can be easily extended to two or three-dimensional systems, which we briefly discuss in Appendix B.
@@
Figure 1: Schematic of layered system in the AGF method. The inhomogeneity or defect, which is represented by the red-shaded oval, is contained within the layers $n=1$ to $N-1$ which correspond to the device region. The distance between the left and right leads is $W=Na$.
@@ II Theoretical basis of AGF-based approach
In the AGF method, (Ong, 2018a, b) which is outlined in Appendix A, we arrange the degrees of freedom in an infinite array of layers, which we enumerate as $n=-\infty,\ldots,0,1,\ldots,\infty$ and assume are evenly spaced with an interlayer distance of $a$, as shown in Fig. 1. Thus, we can associate a discrete coordinate $x=na$ with the center of the $n$-th layer. As before, we partition the system into three subsystems: the left lead ($n\leq 0$), the device ($11<n<N-$ 1), and the right lead ($n\geq N$). In the left and right leads which are homogeneous and identical to one another, each layer acts as a unit cell, has $l$ degrees of freedom and is identical to the other cells in the lead. In the device, the layers are not necessarily homogeneous even though they are spaced evenly and can accommodate the additional degrees of freedom associated with an inhomogeneity. In our derivation, we discuss the scattering cross section associated with the inhomogeneity for an incoming phonon propagating rightward from the left lead towards the device.
@@ II.1 Total outgoing phonon flux from scattering
where $\bm{\Psi}_{\text{inc}}^{\text{L}}(x)=\bm{\Phi}_{\nu k}(x)\equiv\sqrt{|v_{\nu k}|}\bm{u}_{\nu k}(x)e^{ikx}$ represents the incident $\nu k$ phonon mode  [^2] A careful distinction should be made between the terms incident and incoming. The former is used to refer to the unperturbed state when the scatterer is absent while the latter is used to refer to the phonon mode that is traveling towards the device region. A rightward-propagating incident phonon consists of an incoming phonon eigenmode in the left lead and an outgoing phonon eigenmode in the right lead. with group velocity $v_{\nu k}$ and $\bm{u}_{\nu k}(x)$ denotes its Bloch function satisfying the relationship $\bm{u}_{\nu k}(x)=\bm{u}_{\nu k}(x+a)$. Because of flux normalization, Eq. (1) only contains the contribution from the propagating modes and none from the evanescent modes. In the absence of scattering, the phonon wave function at any two layers at $x$ and $x^{\prime}$ only differ by a phase factor determined by its wave vector, i.e., $\bm{\Psi}_{\text{L}}(x)=\bm{\Psi}_{\text{L}}(x^{\prime})e^{ik(x-x^{\prime})}$.
@@
which is written as a sum of the reflected leftward-propagating flux-normalized $\nu^{\prime}k^{\prime}$ eigenmodes associated with the scattering amplitudes $S(\nu^{\prime}k^{\prime},\nu k)$. [^3] The Heaviside term $\Theta(-v_{k^{\prime}})$ eliminates the contributions from the transmitted rightward-propagating waves in Eq. (2). The sum over $\nu^{\prime}k^{\prime}$ in Eq. (2) is over all possible branches and wave vectors at the incident phonon frequency $\omega_{\nu k}$, i.e., $\omega_{\nu^{\prime}k^{\prime}}=\omega_{\nu k}$. We can associate two fluxes with the left-lead wave function: one from the incident mode and the other from the modes reflected from the device into the left lead. The normalized incident flux is $j_{\text{inc}}^{\text{L}}(\nu k)=1$ while the reflected flux is
@@
The RHS of Eq. (4) contains the incident and outgoing forward-scattered wave functions. Each summand has a phase factor of $e^{-ik^{\prime}W}$, where $W$ denotes the width of the scatterer and is taken as the distance between the lead edges (from $n=0$ in the left lead to $n=N$ in the right lead), so we have $W=Na$. In the absence of a scatterer (e.g. homogeneous waveguide), the forward-scattering amplitude in the right lead is $\overline{S}(\nu^{\prime}k^{\prime},\nu k)=\delta_{\nu^{\prime}\nu}\delta_{k^{\prime},k}e^{ik^{\prime}W}$ where $\delta_{k^{\prime},k}$ denotes the Kronecker delta function, and we recover the solution $\overline{\bm{\Psi}}_{\text{R}}(x)=\sum_{\nu^{\prime}k^{\prime}}\overline{S}(\nu^{\prime}k^{\prime},\nu k)\bm{\Phi}_{\nu^{\prime}k^{\prime}}e^{-ik^{\prime}W}\Theta(v_{\nu^{\prime}k^{\prime}})=\bm{\Phi}_{\nu k}(x)$, which is just the incident eigenmode.
@@ II.2 Scattering cross section from the transmission-matrix diagonal element
In the absence of a scatterer, Eq. (9) yields $\sigma=0$ as expected since $S(\nu k,\nu k)=\overline{S}(\nu k,\nu k)$ and $\mathcal{F}(\nu k)=0$. On the other hand, we note that $\sigma\leq 4A$, i.e., the maximum value of the total scattering cross section can be as large as four times the geometrical cross section when $S(\nu k,\nu k)=-\overline{S}(\nu k,\nu k)$ and $\mathcal{F}(\nu k)=-2$(Boya, 2008) While it appears to be counterintuitive and even possibly unphysical, this result also occurs in quantum mechanics for scattering by a hard sphere in three dimensions. (Sakurai and Napolitano, 2020) This is a consequence of our partitioning of the wave function in $\bm{\Psi}_{\text{L}}(x)$ and $\bm{\Psi}_{\text{R}}(x)$ into two non-orthogonal components: the incident wave $\bm{\Psi}_{\text{inc}}$ and the scattered wave $\bm{\Psi}_{\text{scatt}}$ which represents the part of the original wave that has been perturbed. To find $S(\nu k,\nu k)$ for the $\nu k$ mode in the left lead, we use the AGF method to compute the $l\times l$ transmission matrix for the inhomogeneous waveguide, given by (Ong, 2018a, b)
@@
where $\omega=\omega_{\nu k}$ and $\boldsymbol{G}_{\text{RL}}^{\text{ret}}$ denotes the Green’s function between the rightmost left-lead ($n=0$) and the leftmost right-lead ($n=N$) layers. In Eq. (10),$\boldsymbol{V}_{\text{R}}^{\text{ret}}(+)$ and $\boldsymbol{V}_{\text{L}}^{\text{adv}}(-)$ correspond to the velocity matrices for the outgoing right-lead phonons and the incoming left-lead phonons, respectively, while $\boldsymbol{U}_{\text{R}}^{\text{ret}}(+)$ and $\boldsymbol{U}_{\text{L}}^{\text{adv}}(-)$ correspond to the outgoing right-lead eigenmodes and the incoming left-lead eigenmodes, respectively. A more detailed description of these matrices is given in Eqs. (20b)-(22) in Appendix A. The transmission matrix element $[\boldsymbol{t}_{\text{RL}}(\omega)]_{pq}$ connects the flux amplitude of the $p$-th outgoing right-lead modes to the $q$-th incoming left-lead modes for $p,q=1,\ldots,l$. Because the left and right leads are identical, the set of incoming extended left-lead modes, as indexed by $\nu k$, is the same as the set of outgoing extended right-lead modes, and we can order the extended phonon modes by their wave vectors such that $k_{1}<k_{2}<\ldots$.
@@ III.1 Monoatomic harmonic chain with a single isotopic impurity
To illustrate the AGF method for calculating the scattering cross section, we apply it to the simple archetypal example (Allen and Nghiem, 2022) of phonon scattering by a single isotopic impurity of mass $m_{\text{C}}$ in a monoatomic harmonic chain with an interatomic distance of $a$, as shown in Fig. 2(a). In this system, we treat the isotopic impurity as the device; the mass and interatomic force constants of the rest of the system are given by $m$ and $\zeta$, respectively. For the monoatomic harmonic chain, the phonon dispersion is given by $\omega(k)^{2}=\omega_{m}^{2}\sin^{2}(ka/2)$, where $\omega_{m}=2\sqrt{\zeta/m}$, and the phonon group velocity is given by $v(\omega)=\pm\frac{1}{2}a(\omega_{m}^{2}-\omega^{2})^{1/2}$. As there is only one phonon branch, all the terms in Eq. (10) are scalars. Thus, Eq. (10) simplifies to $t_{\text{RL}}(\omega)=i\omega(\omega_{m}^{2}-\omega^{2})^{1/2}G_{\text{RL}}^{\text{ret}}(\omega)$.
@@ III.1.2 Scattering cross section and rate
After some algebra, we obtain from Eq. (12) the expression for the scattering cross section of the single isotopic impurity, normalized by the geometrical cross section $A$,
@@ III.2 Carbon nanotube with an encapsulated C60 molecule
Figure 3: Schematic of the layer arrangement in the (10,10) armchair carbon nanotube with a C60 molecule (pink-shaded atoms) encapsulated within the device region (turquoise-shaded atoms) in layers $n=1$ to $12$. The layers in the leads, which extend semi-infinitely to the left or right, are represented by the yellow-shaded atoms.
@@
Figure 4: (a) Plot of scattering cross section $\sigma/A$ for each rightward-propagating phonon mode (solid circle) in the (10,10) armchair CNT, distributed over the 66 distinct phonon branches. The values of $\sigma/A$ are between 0 and 4 and are indicated by color according to the color bar. (b) The corresponding plot at low frequencies ($\omega<4\times 10^{13}$ rad/s). We observe the three distinct acoustic phonon branches (LA, TW and TA) and the radial breathing mode (BR) as well as the five lowest phonon subbands which are indicated by brown dashed arrows and labeled SB1 to SB5. The phonon branches are also colored according to the percentage of their eigenmode atomic displacements parallel to the longitudinal axis, with the color varying continuously between bright magenta (100%) and black (0%).
@@ III.2.1 Distribution of mode-dependent scattering cross sections
We compute the area-normalized scattering cross section $\sigma(\nu k)/A$ for each rightward-propagating $\nu k$ phonon mode in the (10,10) armchair CNT using Eqs. (10) and (12) for a range of angular frequencies $\omega$ in intervals of $\Delta\omega=2\times 10^{12}$ rad/s from $\omega=2\times 10^{12}$ to $3.16\times 10^{14}$ rad/s. At each $\omega$, we obtain a set of phonon modes with the wave vectors $k_{1},k_{2},\ldots$, which we plot as solid circles in Fig. 4. As in Fig. 2, we indicate the value of $\sigma/A$ of each mode by the color of its solid circle symbol. For ease of interpretation, the solid circle symbols are superimposed on the $\omega-k$ phonon dispersion curves for the (10,10) armchair CNT. (Dresselhaus and Eklund, 2000)
In Fig. 4(a), there are 66 distinct ones due to the degeneracies in the (10,10) armchair CNT. (Dresselhaus and Eklund, 2000) At $k=0$ in the $\omega\rightarrow 0$ limit, there are four acoustic phonon branches corresponding to the longitudinal acoustic (LA), the azimuthal twisting (TW) and the doubly degenerate radial transverse acoustic (TA) phonons. (Mahan, 2002; Chico et al., 2006) In addition, there is a breathing mode (BR) associated with the radial oscillation of the CNT at $k=0$. The remaining phonon branches consist of subbands with $\omega>0$ at $k=0$. We observe two distinct $\sigma/A$ trends for the lower-frequency phonon subbands in Fig. 4(a). In the subbands with higher group velocities and near-linear dispersion, $\sigma/A$ is vanishingly small because the atomic motion for the phonon modes is primarily along the longitudinal direction (Dresselhaus and Eklund, 2000). On the other hand, in the subbands with smaller group velocities and a more parabolic dispersion, $\sigma/A$ increases significantly with $\omega$, indicating stronger interaction with the C60 molecule. For the subbands correspond to the high-frequency optical phonons, we have $\sigma/A\sim 2$, which indicates strong scattering.